The DJ Company issued a 12-year semi-annual pay bond 2 years ago. The face value of the bond is $1,100 and the coupon rate is 4%. The current market rate (annual APR) on comparable bonds is 5%. Calculate the bond's current price.
Answer #1

Number of periods = 10 * 2 = 20

Rate = 5% / 2 = 2.5%

Coupon = (0.04 * 1100) / 2 = 22

Current price = Coupon * [1 -1 / (1 + r)n] / r + FV / (1 + r)n

Current price = 22 * [1 -1 / (1 + 0.025)20] / 0.025 + 1100 / (1 + 0.025)20

Current price = 22 * [1 - 0.610271] / 0.025 + 671.298037

Current price = 22 * 15.58916 + 671.298037

Current price = $1,014.26